If a car starting from rest moves with uniform acceleration and travels the first meter in 1 minute. How long will it take to travel the second meter


1
Expert's answer
2021-03-02T18:05:37-0500
"s=0.5at^2\\1=0.5a(60)^2\\a=\frac{1}{1800}\frac{m}{s^2}\\2=0.5\frac{1}{1800}(t+60)^2\\t=24.85\ s"